# Predicting ionic conductivity in solids from the machine-learned potential energy landscape

### v1

@@ Abstract
Discovering new superionic materials is essential for advancing solid-state batteries, which offer improved energy density and safety compared to the traditional lithium-ion batteries with liquid electrolytes. Conventional computational methods for identifying such materials are resource-intensive and not easily scalable. Recently, universal interatomic potential models have been developed using equivariant graph neural networks. These models are trained on extensive datasets of first-principles force and energy calculations. One can achieve significant computational advantages by leveraging them as the foundation for traditional methods of assessing the ionic conductivity, such as molecular dynamics or nudged elastic band techniques. However, the generalization error from model inference on diverse atomic structures arising in such calculations can compromise the reliability of the results. In this work, we propose an approach for the quick and reliable evaluation of ionic conductivity through the analysis of a universal interatomic potential. Our method incorporates a set of heuristic structure descriptors that effectively employ the rich knowledge of the underlying model while requiring minimal generalization capabilities. Using our descriptors, we rank lithium-containing materials in the Materials Project database according to their expected ionic conductivity. Eight out of the ten highest-ranked materials are confirmed to be superionic at room temperature in first-principles calculations. Notably, our method achieves a speed-up factor of approximately 50 compared to molecular dynamics driven by a machine-learning potential, and is at least 3,000 times faster compared to first-principles molecular dynamics.
@@ I Introduction
To address these challenges, computational prediction methods have been employed to identify and optimize new SSE materials [8, 9, 10, 11, 12]. However, these methods are often computationally expensive and time-consuming. Machine learning (ML) presents a viable solution to mitigate these limitations [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. By leveraging large datasets such as the Materials Project database [25] and advanced algorithms, ML has the potential to speed up the discovery of new materials with high ionic conductivity and therefore accelerate the research and development of high-performance SSBs [26, 27].
Different ML algorithms that have been employed to assist the discovery of superionic materials include classical supervised ML methods like logistic regression [28, 29], gradient boosting [30] and support vector machines [31, 32], classical unsupervised methods [33, 26], as well as deep learning approaches based on graph neural networks (GNN) [34, 35, 36]. The main difficulty in implementing the ML models for identifying SSEs is in general the limited availability of high-quality data on ionic mobility in solids. This motivates the use of knowledge transfer methods [32], or trading off for larger datasets [35] obtained using approximate methods [37].
@@ II Methodology
The overall computational pipeline performed in this work is shown in Fig. 1. We start from a machine-learned interatomic potential, which is used to analyze the potential energy surface (PES) for a given material. Then, a set of scalar heuristics is derived from the obtained PES. These descriptors are based on our intuition of which PES properties should correlate well with ionic mobility. We further validate our choice and evaluate the quality of the derived heuristics with labeled data, both simulated and experimental. Descriptors that demonstrate robustness and highest predictive power are chosen to rank the structures from the Materials Project dataset [25]. Finally, we perform computationally demanding ab initio molecular dynamics (AIMD) simulation to validate the candidates that are most promising according to our method. We describe our pipeline in greater detail throughout Secs. II.1, II.2 and II.3.
@@ III Results and discussion
# MP identifier Composition $\Xi$ AIMD results
$T$ [K] $\tau$ [ps] $\sigma$ [S / cm] $E_{A}$ [meV]
1000 53.698 $2.6\pm 0.6$
1 mp-1211100 LiB3H8 0.9924 667 50.430 $1.18\pm 0.11$ $99\pm 93$
500 52.248 $1.6\pm 0.4$
1000 60.833 $0.9\pm 0.5$
2 mp-1211296 Li(BH)6 0.9898 667 51.756 $0.30\pm 0.15$ $224\pm 19$
500 51.754 $0.13\pm 0.05$
1000 57.168 $3.7\pm 0.4$
3 mp-570935 LiI 0.9812 667 24.413 $-0.016^{*}$ NA
500 66.191 $(2.2\pm 1.0)\times 10^{-3}$
4 mp-721239 Li10Ge(PSe6)2 0.9746 - - - -
5 mp-721252 Li10Sn(PSe6)2 0.9729 - - - -
1000 41.944 $4.19\pm 0.17$
6 mp-2530 Li2Te 0.9712 667 48.011 $(0.46\pm 0.10)\times 10^{-3}$ NA
500 47.926 $-(0.5\pm 1.8)\times 10^{-3}$
7 mp-721253 Li10Si(PSe6)2 0.9712 - - - -
8 mp-705516 Li10Sn(PSe6)2 0.9661 - - - -
1000 65.610 $4.9\pm 0.9$
9 mp-644223 LiBH4 0.9647 667 62.477 $1.4\pm 0.7$ $320\pm 43$
500 72.662 $0.14\pm 0.16$
10 mp-696127 Li10Ge(PSe6)2 0.9645 - - - -
1000 69.483 $1.96\pm 0.14$
12 mp-641703 Li7P3S11 0.9606 667 57.364 $0.8\pm 0.2$ $251\pm 15$
500 59.836 $0.20\pm 0.05$
25 mp-696128 Li10Ge(PS6)2 0.9379 - - - -
Table 1: Top $\Xi$ predictions and AIMD results summary. Ten structures with highest $\Xi$ values are shown. In addition, two known superionic conductors, Li7P3S11 [65] and Li10Ge(PS6)2 (LGPS) [66], are also provided for reference. AIMD simulation runs were conducted with five structures from the top-ten list that are not obtained from LGPS by element substitution [64], and also for one of the reference structures. $E_{A}$ is the activation energy extracted by fitting the Arrhenius law for the diffusivity dependence on $1/T$, where applicable. The * symbol indicates the AIMD simulation run that demonstrated no diffusion within the simulation time interval, and the uncertainty of the extracted diffusion constant is not provided as the MSD fit was only performed in a single time window.
Figure 4: AIMD conductivity values against $1/T$ for the top-$\Xi$ structures. Vertical axis is clipped at minimal conductivity value of $10^{-2}$ S / cm with values below that threshold being indicated as triangles. While the target temperatures in our simulations are exactly 500, 667 and 1000 K, small shifts are added to the horizontal axis values to improve marker visibility. The five LGPS-like structures from the top-$\Xi$ list are omitted from our AIMD studies and therefore not shown in this plot.
@@ III.1 AIMD validation summary
AIMD calculations for the reference Li7P3S11 structure result in the activation barrier estimate of $E_{A}=251\pm 15$ meV and room-temperature-extrapolated conductivity of $7.4\pm 2.9$ mS / cm. These are consistent with the range of values measured experimentally, with $E_{A}=295$ meV and $\sigma=8.3$ mS / cm for the most pure phase sample in Ref. [67] and the lowest activation energy of $E_{A}=124$ meV, with $\sigma=3.2$ mS / cm in Ref. [68].
At the top of the list, we find three hydroborate structures, LiB3H8 (mp-1211100), Li(BH)6 (mp-1211296) and LiBH4 (mp-644223). These belong to the family of lithium hydroborates, of which the best studied is hexagonal LiBH4 [69, 70, 71]. The latter has been found experimentally to have an ionic conductivity of the order of 1 mS / cm at 383 K. Cubic Li(BH)6 has been found to have low diffusivity by a combined DFT and ML-IAP study, but a closely connected orthorhombic distortion of the structure was found to have a conductivity of up to 0.1 S / cm at 700 K [72]. In our AIMD calculations, all the three hydroborates show significant ionic mobility in the studied temperature range, although due to the limited statistics we were unable to reliably extrapolate it to room temperature. Since these hydroborates are molecular solids, the anions have rotational degrees of freedom, and these have been shown to be necessary for the high cationic conductivity in Li(BH)6 [73, 74]. Anionic motion is also apparent on the MSD of the non-Li ions (which we will name the framework) in the AIMD calculations for Li(BH)6 and LiBH4.
LiI (mp-570935) also exhibits high mobility at 1000 K, but accompanied by observable diffusivity of the framework ions, which may indicate that 1000 K is close to the melting point. At low temperature, no diffusivity can be observed. Experimentally, wurtzite LiI is metastable and has been observed [75] but is very hygroscopic and hard to work with. Most ionic conductivity studies employing LiI use cubic LiI as a minor fraction of the electrolyte mixture [76].
@@ IV Conclusion
Our study presents an effective method for the rapid screening of solid electrolyte candidates using heuristic descriptors derived from interatomic potentials. By applying our proposed pipeline, supported by a universal machine-learned interatomic potential, we screened 1302 compounds from the Materials Project database, identifying several promising SSE candidates. AIMD simulations of the 10 highest-ranking materials confirm that all are conductive at high temperatures, with 8 out of 10 being superionic at room temperature. Notably, our approach highlights the solid electrolyte LiB3H8 (mp-1211100), with an impressive AIMD ionic conductivity estimate of $1.6\pm 0.4$ S / cm at $T=500$ K. To our knowledge, this material has not been studied experimentally in the context of SSBs.
@@ VI Acknowledgements
This research project is supported by the Ministry of Education, Singapore, under its Research Centre of Excellence award to the Institute for Functional Intelligent Materials, National University of Singapore (I-FIM, project No. EDUNC-33-18-279-V12). This work used computational resources of the Singapore National Supercomputing Centre (NSCC) of Singapore. For the exploratory phase, this work used computational resources of the Constructor Research Platform provided by Constructor Technologies.
@@ Appendix A Validation with ML-driven MD
MP identifier Composition $\Xi$
mp-1211100 LiB3H8 0.992
mp-1211296 Li(BH)6 0.99
mp-985583 Li3PS4 0.958
mp-1001069 Li48P16S61 0.956
mp-1097034 Li20Si3P3S23Cl 0.951
mp-1040451 Li20Si3P3S23Cl 0.949
mp-1185263 LiPaO3 0.942
mp-1097036 Li3PS4 0.931
mp-1222398 LiGa(GeSe3)2 0.896
mp-755463 Li3SbS3 0.887
mp-753720 Li3BiS3 0.858
mp-1222392 LiUI6 0.82
mp-34038 Li6NCl3 0.814
mp-680395 Li3As7 0.783
mp-775806 Li3SbS3 0.763
mp-28336 Li3P7 0.721
mp-1222582 Li4GeS4 0.717
mp-753429 Li4Bi2S7 0.716
mp-985582 Li6PS5I 0.703
mp-1177520 Li3SbS3 0.658
mp-1211362 Li(BH)6 0.647
mp-1211446 Li7PSe6 0.606
mp-1222482 Li6AsS5I 0.539
mp-1211324 Li7PS6 0.538
mp-1195718 Li4SnS4 0.525
mp-950995 Li6PS5I 0.521
mp-1211176 Li6AsS5I 0.483

### v2

@@ Abstract
Discovering new superionic materials is essential for advancing solid-state batteries, which offer improved energy density and safety compared to the traditional lithium-ion batteries with liquid electrolytes. Conventional computational methods for identifying such materials are resource-intensive and not easily scalable. Recently, universal interatomic potential models have been developed using equivariant graph neural networks. These models are trained on extensive datasets of first-principles force and energy calculations. One can achieve significant computational advantages by leveraging them as the foundation for traditional methods of assessing the ionic conductivity, such as molecular dynamics or nudged elastic band techniques. However, the generalization error from model inference on diverse atomic structures arising in such calculations can compromise the reliability of the results. In this work, we propose an approach for the quick and reliable screening of ionic conductors through the analysis of a universal interatomic potential. Our method incorporates a set of heuristic structure descriptors that effectively employ the rich knowledge of the underlying model while requiring minimal generalization capabilities. Using our descriptors, we rank lithium-containing materials in the Materials Project database according to their expected ionic conductivity. Eight out of the ten highest-ranked materials are confirmed to be superionic at room temperature in first-principles calculations. Notably, our method achieves a speed-up factor of approximately 50 compared to molecular dynamics driven by a machine-learning potential, and is at least 3,000 times faster compared to first-principles molecular dynamics.
@@ I Introduction
To address these challenges, computational prediction methods have been employed to identify and optimize new SSE materials [8, 9, 10, 11, 12]. However, these methods are often computationally expensive and time-consuming. Machine learning (ML) presents a viable solution to mitigate these limitations [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. By leveraging large datasets such as the Materials Project database [26] and advanced algorithms, ML has the potential to speed up the discovery of new materials with high ionic conductivity and therefore accelerate the research and development of high-performance SSBs [27, 28].
Different ML algorithms that have been employed to assist the discovery of superionic materials include classical supervised ML methods like logistic regression [29, 30], gradient boosting [31] and support vector machines [32, 33], classical unsupervised methods [34, 27], as well as deep learning approaches based on graph neural networks (GNN) [35, 36, 37] and transformers [38]. The main difficulty in implementing the ML models for identifying SSEs is in general the limited availability of high-quality data on ionic mobility in solids. This motivates the use of knowledge transfer methods [33], or trading off for larger datasets [36] obtained using approximate methods [39].
@@ II Methodology
The overall computational pipeline performed in this work is shown in Fig. 1. We start from a machine-learned interatomic potential, which is used to analyze the potential energy surface (PES) for a given material. Then, a set of scalar heuristics is derived from the obtained PES. These descriptors are based on our intuition of which PES properties should correlate well with ionic mobility, and are inspired by the frozen host lattice idea behind the pinball model [52] and are similar to the pathway volume assessed with the bond-valence technique in [53]. We further validate our choice and evaluate the quality of the derived heuristics with labeled data, both simulated and experimental. Descriptors that demonstrate robustness and highest predictive power are chosen to rank the structures from the Materials Project dataset [26]. Finally, we perform computationally demanding ab initio molecular dynamics (AIMD) simulation to validate the candidates that are most promising according to our method. We describe our pipeline in greater detail throughout Secs. II.1, II.2 and II.3.
@@ III Results and discussion
# MP identifier [26] Composition $\Xi$ AIMD results
$T$ [K] $\tau$ [ps] $\sigma$ [S / cm] $E_{A}$ [meV] $p$-value ($\frac{\chi^{2}}{df}$)
1 mp-1211100 LiB3H8 0.9924 1000 53.7 $2.6\pm 0.6$ $80\pm 42$ 0.09 (2.4)
667 80.0 $1.27\pm 0.09$
500 104.8 $1.5\pm 0.4$
417 68.9 $1.03\pm 0.30$
363 97.4 $0.082\pm 0.023^{*}$
2 mp-1211296 Li(BH)6 0.9898 1000 60.8 $0.9\pm 0.5$ $219\pm 86$ 0.17 (1.8)
667 81.0 $0.32\pm 0.13$
500 70.3 $0.10\pm 0.05$
417 58.4 $0.26\pm 0.12$
3 mp-570935 LiI 0.9812 1000 57.2 $3.7\pm 0.4$
667 64.4 $0.5\pm 0.6$
500 66.2 $(2.2\pm 1.0)\times 10^{-3}$
4 mp-721239 Li10Ge(PSe6)2 0.9746 - - - - -
5 mp-721252 Li10Sn(PSe6)2 0.9729 - - - - -
6 mp-2530 Li2Te 0.9712 1000 41.9 $4.19\pm 0.17$
667 88.0 $(1.4\pm 0.6)\times 10^{-3}$
500 47.9 $-(0.5\pm 1.8)\times 10^{-3}$
7 mp-721253 Li10Si(PSe6)2 0.9712 - - - - -
8 mp-705516 Li10Sn(PSe6)2 0.9661 - - - - -
9 mp-644223 LiBH4 0.9647 1000 65.6 $4.9\pm 0.9$ $389\pm 26$ 0.54 (0.6)
667 92.9 $1.1\pm 0.4$
500 132.7 $0.24\pm 0.17$
417 92.8 $0.016\pm 0.012$
10 mp-696127 Li10Ge(PSe6)2 0.9645 - - - - -
12 mp-641703 Li7P3S11 0.9606 1000 69.5 $1.96\pm 0.14$ $249\pm 18$ 0.23 (1.5)
667 85.6 $0.86\pm 0.14$
500 119.7 $0.193\pm 0.039$
417 58.1 $0.18\pm 0.09$
25 mp-696128 Li10Ge(PS6)2 0.9379 - - - - -
Table 1: Top $\Xi$ predictions and AIMD results summary. Ten structures with highest $\Xi$ values are shown. In addition, two known superionic conductors, Li7P3S11 [69] and Li10Ge(PS6)2 (LGPS) [70], are also provided for reference. AIMD simulation runs were conducted with five structures from the top-ten list that are not obtained from LGPS by element substitution [68], and also for one of the reference structures. $T$, $\tau$ and $\sigma$ represent the temperature, simulated time, and extracted conductivity value, respectively, for each AIMD simulation run. $E_{A}$ is the activation energy extracted by fitting the Arrhenius law for the diffusivity dependence on $1/T$ using weighted least squares, where applicable. To quantify the goodness of these fits, the $\chi^{2}$-statistic divided by the number of degrees of freedom ($\frac{\chi^{2}}{df}$) is reported along with the corresponding $p$-value. The conductivity measurement marked by the * symbol was excluded from the Arrhenius law fit (see Fig. 4).
Figure 4: AIMD conductivity values against $1/T$ for the top-$\Xi$ structures. Vertical axis is clipped at minimal conductivity value of $10^{-2}$ S / cm with values below that threshold being indicated as triangles. While the target temperatures in our simulations are exactly 500, 667 and 1000 K (with the additional simulations at 417 and 363 K for some materials), small shifts are added to the horizontal axis values to improve marker visibility. The five LGPS-like structures from the top-$\Xi$ list are omitted from our AIMD studies and therefore not shown in this plot.
@@ III.1 AIMD validation summary
AIMD calculations for the reference Li7P3S11 structure result in the activation barrier estimate of $E_{A}=249\pm 18$ meV and room-temperature-extrapolated conductivity of $7.9\pm 3.7$ mS / cm. These are consistent with the range of values measured experimentally, with $E_{A}=295$ meV and $\sigma=8.3$ mS / cm for the most pure phase sample in Ref. [71] and the lowest activation energy of $E_{A}=124$ meV, with $\sigma=3.2$ mS / cm in Ref. [72].
At the top of the list, we find three hydroborate structures, LiB3H8 (mp-1211100), Li(BH)6 (mp-1211296) and LiBH4 (mp-644223). These belong to the family of lithium hydroborates, of which the best studied is hexagonal LiBH4 [73, 74, 75]. The latter has been found experimentally to have an ionic conductivity of the order of 1 mS / cm at 383 K. Cubic Li(BH)6 has been found to have low diffusivity by a combined DFT and ML-IAP study, but a closely connected orthorhombic distortion of the structure was found to have a conductivity of up to 0.1 S / cm at 700 K [76]. In our AIMD calculations, all the three hydroborates show significant ionic mobility in the studied temperature range. For LiBH4, the room-temperature-extrapolated conductivity is $0.44\pm 0.30$ mS / cm. The other two hydroborates deviate from Arrhenius behavior at lower temperatures, preventing reliable extrapolation of their conductivity values to room temperature. Exclusively for LiB3H8, which shows very high conductivity and lowest activation energy in the temperature range $417-1000$ K, we perform additional AIMD simulation at $T=363$ K yielding a conductivity value of $82\pm 23$ mS / cm. This additional measurement is not included in the Arrhenius law fit reported in Table 1 and Fig. 4. Including this point yields $E_{A}=216\pm 31$ meV with a very low $p$-value ($\frac{\chi^{2}}{df}$) of 0.003 (4.7). Since these hydroborates are molecular solids, the anions have rotational degrees of freedom, and these have been shown to be necessary for the high cationic conductivity in Li(BH)6 [77, 78]. Anionic motion is also apparent on the MSD of the non-Li ions (framework) in the AIMD calculations for Li(BH)6 and LiBH4.
LiI (mp-570935) also exhibits high mobility at 1000 K, but accompanied by observable diffusivity of the framework ions, which may indicate that 1000 K is close to the melting point. At $T=500$ K, no diffusivity is observed. Interestingly, in the 667 K run, we see an abrupt change in lithium conductivity from zero during the first 35 ps of the simulation to over 1 S / cm in the subsequent 30 ps, indicating a transition to a highly diffusive state. The value of $0.5\pm 0.6$ S / cm, shown in Table 1 and Fig. 4, is obtained by applying our MSD analysis procedure described in Sec. II.3 to the entire run, which includes both low- and high-diffusivity segments. Experimentally, wurtzite LiI is metastable and has been observed [79] but is very hygroscopic and hard to work with. Most ionic conductivity studies employing LiI use cubic LiI as a minor fraction of the electrolyte mixture [80].
@@ III.3 Discussion
We acknowledge that our descriptors cannot capture the full complexity of the dynamical effects that define the ionic conductivity in solids [85, 86, 87]. However, the effectiveness observed in our screening suggests potential correlations between the complex dynamics of ionic motion and the characteristics derived from the simplified static picture that we utilize.
@@ IV Conclusion
Our study presents an effective method for the rapid screening of solid electrolyte candidates using heuristic descriptors derived from interatomic potentials. By applying our proposed pipeline, supported by a universal machine-learned interatomic potential, we screened 1302 compounds from the Materials Project database, identifying several promising SSE candidates. AIMD simulations of the 10 highest-ranking materials confirm that all are conductive at high temperatures, with 8 out of 10 being superionic at room temperature. Notably, our approach highlights the solid electrolyte LiB3H8 (mp-1211100), with an impressive AIMD ionic conductivity estimate of $82\pm 23$ mS / cm at $T=363$ K. To our knowledge, this material has not been studied experimentally in the context of SSBs.
@@ VI Acknowledgements
This research project is supported by the Ministry of Education, Singapore, under its Research Centre of Excellence award to the Institute for Functional Intelligent Materials, National University of Singapore (I-FIM, project No. EDUNC-33-18-279-V12). This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-028). This work used computational resources of the Singapore National Supercomputing Centre (NSCC) of Singapore. For the exploratory phase, this work used computational resources of the Constructor Research Platform provided by Constructor Technologies.
@@ Appendix A Validation with ML-driven MD
MP identifier Composition System (space group) $\Xi$
mp-1211100 LiB3H8 Orthorhombic (63) 0.992
mp-1211296 Li(BH)6 Cubic (202) 0.99
mp-985583 Li3PS4 Orthorhombic (62) 0.958
mp-1001069 Li48P16S61 Monoclinic (6) 0.956
mp-1097034 Li20Si3P3S23Cl Monoclinic (6) 0.951
mp-1040451 Li20Si3P3S23Cl Monoclinic (6) 0.949
mp-1185263 LiPaO3 Cubic (221) 0.942
mp-1097036 Li3PS4 Orthorhombic (62) 0.931
mp-1222398 LiGa(GeSe3)2 Monoclinic (9) 0.896
mp-755463 Li3SbS3 Trigonal (148) 0.887
mp-753720 Li3BiS3 Trigonal (148) 0.858
mp-1222392 LiUI6 Monoclinic (5) 0.82
mp-34038 Li6NCl3 Monoclinic (8) 0.814
mp-680395 Li3As7 Orthorhombic (61) 0.783
mp-775806 Li3SbS3 Monoclinic (14) 0.763
mp-28336 Li3P7 Orthorhombic (19) 0.721
mp-1222582 Li4GeS4 Orthorhombic (33) 0.717
mp-753429 Li4Bi2S7 Monoclinic (14) 0.716
mp-985582 Li6PS5I Cubic (216) 0.703
mp-1177520 Li3SbS3 Trigonal (160) 0.658
mp-1211362 Li(BH)6 Monoclinic (14) 0.647
mp-1211446 Li7PSe6 Orthorhombic (33) 0.606
mp-1222482 Li6AsS5I Triclinic (1) 0.539
mp-1211324 Li7PS6 Orthorhombic (33) 0.538
mp-1195718 Li4SnS4 Orthorhombic (62) 0.525
mp-950995 Li6PS5I Monoclinic (9) 0.521
mp-1211176 Li6AsS5I Monoclinic (9) 0.483